OpenStudy (anonymous):

Prove it (4-4)/(4-4)=2

OpenStudy (dan815):

lol

OpenStudy (dan815):

given statement is false

OpenStudy (dan815):

as 0/0 is indeterminat or 1/1 = 1 is not equal to 2

OpenStudy (vishweshshrimali5):

As @dan815 mentioned, according to the basic rules of mathematics, dividing by 0 is not allowed and thus you can't cancel out 4-4 from numerator and denominator

OpenStudy (dan815):

sry there is no1/1 its just not possible, its just 0/0 is indeterminate

OpenStudy (dan815):

yeah i just thought about it :P

OpenStudy (vishweshshrimali5):

Actually 0/0 form is very interesting. According to some people, it is indeterminate because it can have different "finite" values depending on different cases

OpenStudy (dan815):

because to cancel 4-4 you have to divide top and bottom by 4-4 in which case yo run into your previous indeterminatet problem

OpenStudy (vishweshshrimali5):

So, by some mathematical trick you "can" show that 4-4/4-4 = 2 (this sort of things are very common in 0/0 formats)

OpenStudy (vishweshshrimali5):

Though, that particular mathematical trick would have some mistake for sure

OpenStudy (dan815):

yes

OpenStudy (dan815):

maybe you can arrive at any answer if you discard division by 0

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 For example: \[\large{\cfrac{4-4}{4-4} = \cfrac{4(1-1)}{4(1-1)} = \cfrac{1-1}{1-1} = \cfrac{1^2 - 1^2}{1-1} = \cfrac{(1-1)(1+1)}{1-1} = 1+1}\] \[\large{ = 2}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

^ these are called mathematical fallacies

OpenStudy (dan815):



OpenStudy (dan815):

u can simpy multiply rightside by 4-4 and add 8 to both sides

OpenStudy (dan815):